# Contextualized Counterspeech: Strategies for Adaptation, Personalization, and Evaluation

### v2

@@ Abstract.
AI-generated counterspeech offers a promising and scalable strategy to curb online toxicity through direct replies that promote civil discourse. However, current counterspeech is one-size-fits-all, lacking adaptation to the moderation context and the users involved. We propose and evaluate multiple strategies for generating tailored counterspeech that is adapted to the moderation context and personalized for the moderated user. We instruct an LLaMA2-13B model to generate counterspeech, experimenting with various configurations based on different contextual information and fine-tuning strategies. We identify the configurations that generate persuasive counterspeech through a combination of quantitative indicators and human evaluations collected via a pre-registered mixed-design crowdsourcing experiment. Results show that contextualized counterspeech can significantly outperform state-of-the-art generic counterspeech in adequacy and persuasiveness, without compromising other characteristics. Our findings also reveal a poor correlation between quantitative indicators and human evaluations, suggesting that these methods assess different aspects and highlighting the need for nuanced evaluation methodologies. The effectiveness of contextualized AI-generated counterspeech and the divergence between human and algorithmic evaluations underscore the importance of increased human-AI collaboration in content moderation.
@@ 1. Introduction
Platforms apply a wide array of moderation actions to curb toxicity and other online harms (Trujillo et al., 2025; Dwork et al., 2024). These include content and user removals, friction interventions, and demotion (Trujillo and Cresci, 2022; Chandrasekharan et al., 2022). An alternative to centralized moderation is social correction—such as counterspeech—where users proactively respond to toxic content encouraging respectful and constructive communication (Garland et al., 2022; Bonaldi et al., 2024). Counterspeech holds promise as it avoids censorship concerns while preventing toxic users from merely relocating to other communities (Horta Ribeiro et al., 2021). However, a number of challenges currently limit its efficacy and applicability. The emotional toll on responders, who spend countless hours confronting hate, incivility, and personal attacks, is a major issue (Roberts, 2019; Steiger et al., 2021). Safety is another concern, as responders may face retaliation (Doerfler et al., 2021; Tabassum et al., 2024). Moreover, the widespread nature of online toxicity makes manual counterspeech highly impractical. To overcome these limitations, automated counterspeech systems leveraging generative AI technologies—such as large language models (LLMs)—have been developed (Tekiroğlu et al., 2020; Bonaldi et al., 2024). Yet, evaluating their effectiveness remains challenging (Gillespie, 2020). Existing work primarily focused on basic characteristics like grammaticality and relevance, using quantitative indicators that offer a narrow assessment of effectiveness (Zubiaga et al., 2024). Additionally, current machine-generated counterspeech is one-size-fits-all (Cresci et al., 2022), relying solely on the toxic message as input, which is often insufficient for crafting effective responses. On the contrary, effective moderation must be context-dependent (Gillespie, 2020; Huang et al., 2024), requiring a shift from generic to tailored approaches that consider the broader conversational context (Cresci et al., 2022).
@@ 2.1. LLM-generated counterspeech
Counterspeech can be more effective than other moderation actions, such as content and user removals, without limiting free speech (Chung et al., 2023). Favorable results were obtained by scholars (Hangartner et al., 2021; Bilewicz et al., 2021), NGOs (Chung et al., 2019, 2021), and ordinary users (Goffredo et al., 2022; Hassan and Alikhani, 2023) alike, as part of observational (Garland et al., 2022), quasi-experimental (Bilewicz et al., 2021), and experimental (Munger, 2017; Hangartner et al., 2021) studies. These positive results also extend to AI-generated counterspeech. However, we currently lack a thorough understanding of the types of counterspeech that are most effective and the optimal conditions for their generation (Chung et al., 2023). For these reasons, research is focusing on identifying the conditions, methods, and resources that make AI-generated counterspeech effective. Tekiroglu et al. (2022) compared pre-trained LLMs and decoding strategies to highlight those capable of generating effective counterspeech. Others leveraged socio-psychological theories to generate messages based on different strategies, such as empathy, abstract norms, disapproval, and humor (Sue et al., 2019; Hangartner et al., 2021; Bilewicz et al., 2021; Bär et al., 2024). However, these strategies were applied independently of the moderated context and user. As shown in Table 1, the only exceptions are Doğanç and Markov (2023) who used the age and gender of hateful users to generate tailored counterspeech, and Bär et al. (2024) who prompted an LLM to generate contextualized counterspeech based on the content of the toxic message. Herein, we go beyond existing work by generating adapted and personalized counterspeech that leverage content about the user, the conversation, and the community, assessing which type of contextual information results in effective counterspeech.
@@ 2.2. Persuasiveness of LLM-generated messages
LLMs persuasiveness has been explored from various perspectives. While some studies focused on persuasion directed at other LLMs (Breum et al., 2024; Salvi et al., 2024), most research examined their influence on humans. For instance,  (Furumai et al., 2024; Shi et al., 2020) explored attempts to secure donations through different communication strategies, while (Ramani et al., 2024) used LLMs as sales agents. Propaganda was the focus of  (Goldstein et al., 2024), which demonstrated that AI-generated and human-generated propaganda exhibit comparable levels of persuasiveness, suggesting they can be used interchangeably. Similar results were obtained about political persuasiveness (Hackenburg et al., 2024). Costello et al. (2024) instructed an LLM to debunk conspiracy theories, achieving notable long-term effects, with persuasion lasting up to two months post-intervention. Finally, several studies examined the use of LLMs as a moderation tool for mitigating conflicts and for reducing toxic messages on social media(Cho et al., 2023; Hong et al., 2024; Govers et al., 2024). In spite of these results however, the persuasiveness of AI-generated contextualized counterspeech is still unexplored.
@@ 2.3. LLM adaptation and personalization
The majority of existing content moderation interventions is one-size-fits-all, where the intervention is applied uniformly to all moderated users. However, multiple studies have highlighted the limitations of this generic strategy (Trujillo and Cresci, 2023; Costello et al., 2024; Cima et al., 2024a), suggesting that contextualized moderation could offer substantial improvements (Cresci et al., 2022). However, despite the potential benefits, contextualized moderation has received little attention. Among the few existing works that used LLMs to generate tailored moderation interventions is (Costello et al., 2024), which experimented with personalized messages to debunk conspiracy beliefs. Similarly, Bär et al. (2024) adapted counterspeech to align with the specific content of the moderated messages. Besides content moderation, LLM personalization was studied for socio-demographic alignment (Salvi et al., 2024; Doğanç and Markov, 2023; Beck et al., 2024; Giorgi et al., 2024), for alignment with specific personality traits (Jiang et al., 2023; Shi et al., 2020), and for political microtargeting (Hackenburg and Margetts, 2024), with promising results. Herein, we extend this line of research by experimenting with multiple strategies of adaptation and personalization to generate persuasive contextualized counterspeech.
@@ 3. Problem definition
Let $\mathbf{T}=\langle m_{0},m_{1},\ldots,m_{N}\rangle$ be an online conversation thread where $m_{0},m_{1},\ldots,m_{N}$ denotes the chronologically ordered sequence of thread messages. Given a toxic message $m_{i}$, we aim to generate the counterspeech message $\hat{m}_{i+1}=\mathcal{G}(m_{i},\mathbf{C}_{i})$, where $\mathcal{G}$ is the counterspeech generator and $\mathbf{C}_{i}$ is the contextual information. Contrarily to existing works where $\mathcal{G}$ only receives $m_{i}$ as input (Cho et al., 2023; He et al., 2023; Hong et al., 2024; Leekha et al., 2024; Bär et al., 2024), we generate adapted and personalized counterspeech by also providing $\mathbf{C}_{i}$ as input to $\mathcal{G}$.
Relevant properties. We identify the following set of desired properties that effective counterspeech should possess:
In addition to the above, we also consider the following relevant properties of contextualized AI-generated counterspeech:
We do not consider basic linguistic properties such as fluency or grammaticality, since modern LLMs are capable of consistently generating human-quality text (Li et al., 2024).
@@ 4.1. Generation
We use an instruction-tuned version of LLaMA2-13B to generate counterspeech responses (Touvron et al., 2023). [^1] https://huggingface.co/dfurman/Llama-2-13B-Instruct-v0.2 The instruction prompt is reported in Appendix A. Starting from this generator, we evaluate different configurations depending on the information provided to the model and the data used for fine-tuning. Below, we describe the factors considered in our experiments, each identified by a unique [label]. Some factors do not involve adaptation nor personalization:
@@ 4.1.2. Personalization
  • Summary [Su]: Given a toxic message $m_{i}$, we feed twenty previous messages of its author to an instruction-tuned LLaMA2-13B model tasked with producing user summaries that highlight writing style, lexicon, and main interests. The instruction prompt used to generate user summaries is reported in Appendix A. User summaries are then provided to the counterspeech generator as a source of user-specific information.
@@ 4.2.3. Human evaluation
We carry out an extensive human evaluation campaign via a pre-registered, [^4] https://aspredicted.org/b55z-5qy4.pdf mixed design crowdsourcing experiment on Amazon Mechanical Turk. [^5] We received ethical clearance (#0306210) for this experiment from the competent IRB. Initially, participants are assigned to one of two between-subjects conditions: (i) non-contextual – participants are shown only the toxic messages and the corresponding generated counterspeech responses; (ii) contextual – participants are also shown the contextual information that was used to adapt and/or personalize the counterspeech. After this initial assignment, participants are asked to read and evaluate multiple pairs ($m_{i},\hat{m}_{i+1}$) of toxic messages and counterspeech, under seven within-subjects conditions. The seven conditions correspond to the seven configurations selected for human evaluation. The order of within-subjects conditions is randomized to control for order effects. In both within-subjects experiments participants are asked to rate on a five points Likert scale: the relevance of the counterspeech to the toxic message, its adequacy as counterspeech, its truthfulness, its artificiality, and its persuasiveness. Following recent literature (Hong et al., 2024), persuasiveness is evaluated in two different questions based on the counterspeech likelihood of (i) persuading the author of the toxic message to re-engage in the conversation in a civil manner, and of (ii) steering the conversation back to civil discourse. Participants assigned to the contextual between-subjects condition are also asked to rate how contextualized is the counterspeech to the context of the toxic message. Finally, all participants are asked a small set of socio-demographic questions. The complete list of questions is reported in Appendix B. This mixed design allows us to isolate and evaluate the impact of contextual information on the effectiveness of counterspeech. In particular, the between-subjects design enables us to determine whether the inclusion of contextual information enhances the perceived relevance, adequacy, and persuasiveness of the responses.
@@ 5. Data
We collected Reddit comments posted over multiple years in five popular subreddits discussing US politics. We selected subreddits with a marked right- (r/conservatives) or left-leaning (r/progressive) tendency, with a focus on a prominent conservative (r/the_donald for Donald Trump) or progressive (r/aoc for Alexandra Ocasio-Cortez) figure, and a mixed-leaning subreddit (r/politics). Depending on subreddit size, data collection spanned either 36 or 12 months so as to collect a comparable amount of data from each subreddit. For r/the_donald, data collection stopped in June 2020, when the subreddit was permanently banned (Cima et al., 2024b). All data was collected from the Pushshift archives (Baumgartner et al., 2020).
@@
Adaptation and personalization datasets. We also built a few additional datasets that implement the adaptation and personalization strategies described in Section 4.1. We selected a stratified random sample of around 7,500 comment-reply pairs that we used to fine-tune the counterspeech generator to Reddit’s political conversational style and informal language, thus implementing the community adaptation factor [Re]. We also selected the two preceding comments to each of the 128 previously selected toxic comments, which we fed to the counterspeech generator via prompting, implementing the conversation adaptation factor [Pr]. Finally, we collected twenty random comments for each author of the 128 selected toxic comments. These data were used to generate the user profiles that implement the user summary personalization factor [Su]. Ten of these comments were also directly fed via prompting to the counterspeech generator, implementing the comment history factor [Hi].
@@ 6.2. Human evaluation
We recruited $N=2,444$ and $N=2,353$ participants on Amazon Mechanical Turk, respectively for the non-contextual and contextual between-subjects experimental conditions. These figures exclude participants whose answers were rejected due to excessively fast survey completion time, $<100\%$ correct answers to the control questions, and those that gave the same answer to all questions. We report no deviations from the pre-registration protocol.
@@ 6.2.2. Contextual experiment
Participants assigned to the contextual condition where shown additional information on top of the toxic message and the generated counterspeech. Contextual information includes the name of the subreddit where the toxic message was posted, the previous message in the conversation thread, and the user summary obtained as described in Section 4.1.2. Again, the Friedman test reveals statistically significant differences. Detailed results are presented in Figure 4 and largely corroborate those from the non-contextual experiment. The configurations [BaPr] and [BaPrHi] consistently achieve the highest overall scores, with [BaPr] demonstrating a statistically significant improvement over the baseline in persuading the author of the toxic message. Most other results for these configurations are statistically non-significant. Conversely, configurations [MuRe], [HsHi], [MuHsHi], and [MuRePrHi] consistently perform significantly worse than the baseline across all metrics, except for [MuRe] and [HsHi] who outperform the baseline in terms of artificiality. This experiment contains an additional question with respect to the non-contextual one, where participants rated the effectiveness of contextualization. Figure 4G shows that while most differences are non-significant, [MuRePrHi] performs worse and [BaPrHi] performs markedly better than the baseline, which reinforces previous results about the effectiveness of the [BaPrHi] configuration.

### v3

@@ Abstract.
AI-generated counterspeech offers a promising and scalable strategy to curb online toxicity through direct replies that promote civil discourse. However, current counterspeech is one-size-fits-all, lacking adaptation to the moderation context and the users involved. We propose and evaluate multiple strategies for generating tailored counterspeech that is adapted to the moderation context and personalized for the moderated user. We instruct a LLaMA2-13B model to generate counterspeech, experimenting with various configurations based on different contextual information and fine-tuning strategies. We identify the configurations that generate persuasive counterspeech through a combination of quantitative indicators and human evaluations collected via a pre-registered mixed-design crowdsourcing experiment. Results show that contextualized counterspeech can significantly outperform state-of-the-art generic counterspeech in adequacy and persuasiveness, without compromising other characteristics. Our findings also reveal a poor correlation between quantitative indicators and human evaluations, suggesting that these methods assess different aspects and highlighting the need for nuanced evaluation methodologies. The effectiveness of contextualized AI-generated counterspeech and the divergence between human and algorithmic evaluations underscore the importance of increased human-AI collaboration in content moderation.
@@
Counterspeech; generative AI; personalization; content moderation; online toxicity
copyright: noneconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714507isbn: 979-8-4007-1274-6/25/04ccs: Computing methodologies Natural language generationccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
Platforms apply a wide array of moderation actions to curb toxicity and other online harms (Trujillo et al., 2025; Dwork et al., 2024). These include content and user removals, friction interventions, and demotion (Trujillo and Cresci, 2022; Chandrasekharan et al., 2022; Tessa et al., 2024). An alternative to centralized moderation is social correction—such as counterspeech—where users proactively respond to toxic content encouraging respectful and constructive communication (Garland et al., 2022; Bonaldi et al., 2024). Counterspeech holds promise as it avoids censorship concerns while preventing toxic users from merely relocating to other communities (Horta Ribeiro et al., 2021). However, a number of challenges currently limit its efficacy and applicability. The emotional toll on responders, who spend countless hours confronting hate, incivility, and personal attacks, is a major issue (Roberts, 2019; Steiger et al., 2021). Safety is another concern, as responders may face retaliation (Doerfler et al., 2021; Tabassum et al., 2024). Moreover, the widespread nature of online toxicity makes manual counterspeech highly impractical. To overcome these limitations, automated counterspeech systems leveraging generative AI technologies—such as large language models (LLMs)—have been developed (Tekiroğlu et al., 2020; Bonaldi et al., 2024). Yet, evaluating their effectiveness remains challenging (Gillespie, 2020). Existing work primarily focused on basic characteristics like grammaticality and relevance, using quantitative indicators that offer a narrow assessment of effectiveness (Zubiaga et al., 2024). Additionally, current machine-generated counterspeech is one-size-fits-all (Cresci et al., 2022), relying solely on the toxic message as input, which is often insufficient for crafting effective responses. On the contrary, effective moderation must be context-dependent (Gillespie, 2020; Huang et al., 2024), requiring a shift from generic to tailored approaches that consider the broader conversational context (Cresci et al., 2022).
@@ 2.1. LLM-generated counterspeech
Counterspeech can be more effective than other moderation actions, such as content and user removals, without limiting free speech (Chung et al., 2023). Favorable results were obtained by scholars (Hangartner et al., 2021; Bilewicz et al., 2021), NGOs (Chung et al., 2019, 2021), and ordinary users (Goffredo et al., 2022; Hassan and Alikhani, 2023) alike, as part of observational (Garland et al., 2022), quasi-experimental (Bilewicz et al., 2021), and experimental (Munger, 2017; Hangartner et al., 2021) studies. These positive results also extend to AI-generated counterspeech. However, we currently lack a thorough understanding of the types of counterspeech that are most effective and the optimal conditions for their generation (Chung et al., 2023). For these reasons, research is focusing on identifying the conditions, methods, and resources that make AI-generated counterspeech effective. Tekiroglu et al. (2022) compared pre-trained LLMs and decoding strategies to highlight those capable of generating effective counterspeech. Multiple studies leveraged socio-psychological theories to generate messages based on different strategies, such as empathy, abstract norms, disapproval, and humor (Sue et al., 2019; Hangartner et al., 2021; Bilewicz et al., 2021; Bär et al., 2024). Others experimented with Retrieval-Augmented Generation (RAG) to integrate external knowledge flexibly and improve the specificity of the generated counterspeech (Jiang et al., 2025). However, all these strategies were applied independently of the moderated context and user. As shown in Appendix Table 2, the only exceptions are Doğanç and Markov (2023) who used the age and gender of hateful users to generate tailored counterspeech, and Bär et al. (2024) who prompted an LLM to generate contextualized counterspeech based on the content of the toxic message. Herein, we go beyond existing work by generating adapted and personalized counterspeech that leverage content about the user, the conversation, and the community, assessing which type of contextual information results in effective counterspeech.
@@ 2.2. Persuasiveness of LLM-generated messages
LLMs persuasiveness has been explored from various perspectives. While some studies focused on persuasion directed at other LLMs (Breum et al., 2024; Salvi et al., 2024), most research examined their influence on humans. For instance, (Furumai et al., 2024; Shi et al., 2020) explored attempts to secure donations through different communication strategies, while (Ramani et al., 2024) used LLMs as sales agents. Propaganda was the focus of (Goldstein et al., 2024), which demonstrated that AI-generated and human-generated propaganda exhibit comparable levels of persuasiveness, suggesting they can be used interchangeably. Similar results were obtained about political persuasiveness (Hackenburg et al., 2024). Instead, Costello et al. (2024) instructed an LLM to debunk conspiracy theories, achieving notable long-term effects, with persuasion lasting up to two months post-intervention. Finally, several studies examined the use of LLMs as a moderation tool for mitigating conflicts and for reducing toxic messages on social media (Cho et al., 2024; Hong et al., 2024; Govers et al., 2024). In spite of these results however, the persuasiveness of AI-generated contextualized counterspeech is still unexplored.
@@ 2.3. LLM adaptation and personalization
The majority of existing content moderation interventions is one-size-fits-all, where the intervention is applied uniformly to all moderated users. However, multiple studies have highlighted the limitations of this generic strategy (Trujillo and Cresci, 2023; Costello et al., 2024; Cima et al., 2024a), suggesting that contextualized moderation could offer substantial improvements (Cresci et al., 2022). However, despite the potential benefits, contextualized moderation has received little attention. Among the few existing works that used LLMs to generate tailored moderation interventions is (Costello et al., 2024), which experimented with personalized messages to debunk conspiracy beliefs. Similarly, Bär et al. (2024) adapted counterspeech to align with the specific content of the moderated messages. Besides content moderation, LLM personalization was studied for socio-demographic alignment (Salvi et al., 2024; Doğanç and Markov, 2023; Beck et al., 2024; Giorgi et al., 2025), for alignment with specific personality traits (Jiang et al., 2024; Shi et al., 2020), and for political microtargeting (Hackenburg and Margetts, 2024), with promising results. Zugecova et al. (2024) also evaluated the effectiveness of personalized disinformation. Herein, we extend this line of research by experimenting with multiple strategies of adaptation and personalization to generate persuasive contextualized counterspeech.
@@ 3. Problem definition
Let $\mathbf{T}=\langle m_{0},m_{1},\ldots,m_{N}\rangle$ be an online conversation thread where $m_{0},m_{1},\ldots,m_{N}$ denotes the chronologically ordered sequence of thread messages. Given a toxic message $m_{i}$, we aim to generate the counterspeech message $\hat{m}_{i+1}=\mathcal{G}(m_{i},\mathbf{C}_{i})$, where $\mathcal{G}$ is the counterspeech generator and $\mathbf{C}_{i}$ is the contextual information. Contrarily to existing works where $\mathcal{G}$ only receives $m_{i}$ as input (Cho et al., 2024; He et al., 2023; Hong et al., 2024; Leekha et al., 2024; Bär et al., 2024), we generate adapted and personalized counterspeech by also providing $\mathbf{C}_{i}$ as input to $\mathcal{G}$.
Properties. We identify the following set of desired properties that effective counterspeech should possess: politeness, adequacy, relevance, diversity, truthfulness, persuasiveness. In addition to the above, we also consider the following relevant properties of contextualized AI-generated counterspeech: adaptation, personalization, artificiality. Each of these properties is defined and motivated in Appendix Section B. Instead, we do not consider basic linguistic properties such as fluency or grammaticality, since modern LLMs are capable of consistently generating human-quality text (Li et al., 2024).
@@ 4.1. Generation
We use an instruction-tuned version of LLaMA2-13B [^1] https://huggingface.co/dfurman/Llama-2-13B-Instruct-v0.2 to generate counterspeech responses (Touvron et al., 2023). The instruction prompts are reported in Appendix C. Starting from this generator, we evaluate different configurations depending on the information provided to the model and the data used for fine-tuning. [^2] Our models are available at https://huggingface.co/collections/alemiaschi/contextualized-counterspeech-llama-2-models-679e322e663f033e1aa654f2 Below, we describe the factors considered in our experiments, each identified by a unique [label]. Some factors do not involve adaptation nor personalization:
@@ 4.1.2. Personalization
  • Summary [Su]: Given a toxic message $m_{i}$, we feed twenty previous messages of its author to an instruction-tuned LLaMA2-13B model tasked with producing user summaries that highlight writing style, lexicon, and main interests. The instruction prompt used to generate user summaries is reported in Appendix C. User summaries are then provided to the counterspeech generator as a source of user-specific information.
@@ 4.2.3. Human evaluation
We carry out an extensive human evaluation campaign via a pre-registered, [^5] https://aspredicted.org/b55z-5qy4.pdf mixed design crowdsourcing experiment on Amazon Mechanical Turk. [^6] We received ethical clearance (#0306210) for this research from CNR’s IRB. Initially, participants are assigned to one of two between-subjects conditions: (i) non-contextual – participants are shown only the toxic messages and the corresponding generated counterspeech responses; (ii) contextual – participants are also shown the contextual information that was used to adapt and/or personalize the counterspeech. After this initial assignment, participants are asked to read and evaluate multiple pairs ($m_{i},\hat{m}_{i+1}$) of toxic messages and counterspeech, under seven within-subjects conditions. The seven conditions correspond to the seven configurations selected for human evaluation. The order of within-subjects conditions is randomized to control for order effects. In both within-subjects experiments participants are asked to rate on a five points Likert scale: the relevance of the counterspeech to the toxic message, its adequacy as counterspeech, its truthfulness, its artificiality, and its persuasiveness. Following recent literature (Hong et al., 2024), persuasiveness is evaluated in two different questions based on the counterspeech likelihood of (i) persuading the author of the toxic message to re-engage in the conversation in a civil manner, and of (ii) steering the conversation back to civil discourse. Participants assigned to the contextual between-subjects condition are also asked to rate how contextualized is the counterspeech to the context of the toxic message. Finally, all participants are asked a small set of socio-demographic questions. The complete list of questions is reported in Appendix D. This mixed design allows us to isolate and evaluate the impact of contextual information on the effectiveness of counterspeech. In particular, the between-subjects design enables us to determine whether the inclusion of contextual information enhances the perceived relevance, adequacy, and persuasiveness of the responses.
@@ 5. Data
We collected Reddit comments posted over multiple years in five popular subreddits discussing US politics. We selected subreddits with a right- (r/conservatives) or left-leaning (r/progressive) tendency, with a focus on a prominent conservative (r/the_donald for Donald Trump) or progressive (r/aoc for Alexandra Ocasio-Cortez) figure, and a mixed-leaning subreddit (r/politics). Depending on subreddit size, data collection spanned either 36 or 12 months so as to collect a comparable amount of data from each subreddit. For r/the_donald, data collection stopped in June 2020, when the subreddit was permanently banned (Cima et al., 2024b). All data was collected from the Pushshift archives (Baumgartner et al., 2020).
@@
Adaptation and personalization datasets. We also built a few additional datasets that implement the adaptation and personalization strategies described in Section 4.1. We selected a stratified random sample of around 7,500 comment-reply pairs that we used to fine-tune the counterspeech generator to Reddit’s political conversational style and language, thus implementing the community adaptation factor [Re]. We also selected the two preceding comments to each of the 128 previously selected toxic comments, which we fed to the counterspeech generator via prompting, implementing the conversation adaptation factor [Pr]. Finally, we collected twenty random comments for each author of the 128 selected toxic comments. These data were used to generate the user profiles that implement the user summary personalization factor [Su]. Ten of these comments were also directly fed via prompting to the counterspeech generator, implementing the comment history factor [Hi].
@@ 6.2. Human evaluation
We recruited $N=2,444$ and $N=2,353$ participants on Amazon Mechanical Turk, respectively for the non-contextual and contextual between-subjects experimental conditions. These figures exclude participants whose answers were rejected due to excessively fast survey completion time, $<100\%$ correct answers to the control questions, and those who gave the same answer to all questions. We report no deviations from the pre-registration protocol.
@@ 6.2.2. Contextual experiment
Participants assigned to the contextual condition where shown additional information on top of the toxic message and the generated counterspeech. Contextual information includes the name of the subreddit where the toxic message was posted, the previous message in the conversation thread, and the user summary obtained as described in Section 4.1.2. Again, the Friedman test reveals statistically significant differences. Detailed results are presented in Figure 4 and largely corroborate those from the non-contextual experiment. The configurations [BaPr] and [BaPrHi] consistently achieve the highest overall scores, with [BaPr] demonstrating a statistically significant improvement over the baseline in persuading the author of the toxic message. Most other results for these configurations are statistically non-significant. Conversely, configurations [MuRe], [HsHi], [MuHsHi], and [MuRePrHi] consistently perform significantly worse than the baseline across all metrics, except for [MuRe] and [HsHi] who outperform the baseline in terms of artificiality. This experiment contains an additional question with respect to the non-contextual one, where participants rated the effectiveness of contextualization. Figure 4G shows that while most differences are non-significant, [MuRePrHi] performs worse and [BaPrHi] performs markedly better than the baseline, which reinforces previous results about the effectiveness of [BaPrHi].